9.1 Find the following values for a single cash flow: a. The future value of $500 invested at 8 percent for one year b. The f draw time line for each question (a,b,c,d)
Answer #1

Ans a) Future value = present value * (1 + interest rate) ^n

= 500 * 1.08

= $540

Ans b) Future value = present value * (1 + interest value)^n

= 500 * (1.08)^5

= $734.66

Ans c) Present value = future value/(1 + interest rate)^n

= 500/1.08

= $462.96

Ans d) Present value = future value/(1 + interest rate)^n

= 500/1.08^5

= $340.29

Time line for a

$500 $540 Present value Future value at year 1

Time line for b

$500 $ 734.66 Present value Future value at year 5

Time line for c

$462.96 $ 500 Present value Future value at year 1

Time line of d

$340.29 $ 500 Present value Future value at year 5